OpenStudy (anonymous):

The volume of a box is given the formula v=lwh volume of the box is 480cu units length given is x+2 width is x. find the width and length.

OpenStudy (mertsj):

What is the height?

OpenStudy (anonymous):

Is any information on the height given?

OpenStudy (mertsj):

x(x+2)(h)=480

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

If h = 6 you can use Mertsj's equation up there and just solve for x

OpenStudy (mertsj):

6(x^2+2x)=480 6x^2+12x-480=0 x^2+2x-80=0 (x+10)(x-8)=0 x=-10 or 8 Reject -10 x=8 and x+2=10

OpenStudy (anonymous):

muahsssssss